Consider the combustion of propane: Assume that al
Answer #1

Consider the combustion of propane:

C3H8(g) + 5O2(g) → 3CO2(g) + 4H2O(l)

ΔH = –2221 kJ
–2221 kJ is the heat given off when one mole of propane is combusted.
Moles of propane = wt/ m.wt
= 5g/44.1g/mol
= 0.1134 moles of propane
ΔH = –2221 kJ x 0.1134
= - 251.86 kj

Soln 2.
50 ml * 0.997 g H2O / ml = 49.85 g H2O

(49.85 g H2O)*(4.180 J / g - C)*(32.3 - 25) = 1562.8 J

MW of CaO = 56.077 g/mol

mol CaO = 1.045 / 56.077 g/mol = 0.0186 mol

1562.8 J / 0.0186 mol = 84000 J / mol = 84 kJ / mol

and it's exothermic so it would be
= -84 kJ/mol